# Chow groups of one-dimensional noetherian domains

### v2

@@ 4. Further examples
The maximal order ${\tilde{\mathcal{O}}}$ of $K:=\mathbb{Q}(\sqrt{-3},\sqrt{13})$ has class number $2$. Lemma 3.4 shows that the order $\mathcal{O}=\mathbb{Z}+2{\tilde{\mathcal{O}}}$ has a unique prime ideal $\mathfrak{p}$ over $2$. Then $2{\tilde{\mathcal{O}}}=\mathfrak{P}\mathfrak{Q}$ with maximal ideals $\mathfrak{P}$ and $\mathfrak{Q}$ which both generate the class group of ${\tilde{\mathcal{O}}}$. Since $\mathfrak{P}$ and $\mathfrak{Q}$ have degree $2$ over $\mathfrak{p}$, the local Chow group $\operatorname{Chow}(\mathcal{O}_{\mathfrak{p}})$ has order $2$. Theorem 2.9 shows
@@
splits into two prime ideals with $[{\tilde{\mathcal{O}}}:\mathfrak{P}_{1}]=7^{2}$ and $[{\tilde{\mathcal{O}}}:\mathfrak{P}_{2}]=7^{3}$. The order
@@
has conductor $\mathfrak{P}_{1}$ and index $7$ in ${\tilde{\mathcal{O}}}$. By Lemma 3.4 there exists a unique prime ideal $\mathfrak{p}_{1}$ of $\mathcal{O}_{1}$ below $\mathfrak{P}_{1}$. Since $\mathcal{O}_{1}/\mathfrak{p}_{1}\cong\mathbb{F}_{7}$, Theorem 2.9 shows that
@@
has conductor $\mathfrak{P}_{2}$ and index $7^{2}$ in ${\tilde{\mathcal{O}}}$. There exists a unique prime ideal $\mathfrak{p}_{2}$ of $\mathcal{O}_{2}$ below $\mathfrak{P}_{2}$. Again, $\mathcal{O}_{1}/\mathfrak{p}_{2}\cong\mathbb{F}_{7}$ implies that
@@
is generated by $[\mathfrak{p}_{2}]$. The intersection $\mathcal{O}^{\prime}:=\mathcal{O}_{1}\cap\mathcal{O}_{2}$ has conductor $7{\tilde{\mathcal{O}}}$ and index $7^{3}$ in ${\tilde{\mathcal{O}}}$. It has two different prime ideals over $7$, namely $\mathfrak{q}_{i}:=\mathcal{O}^{\prime}\cap\mathfrak{p}_{i}$ for $i=1,2$. Theorem 2.9 shows that
$\operatorname{Chow}(\mathcal{O}^{\prime})\cong\operatorname{Chow}(\mathcal{O}^{\prime}_{\mathfrak{q}_{1}})\times\operatorname{Chow}(\mathcal{O}^{\prime}_{\mathfrak{q}_{1}})\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z}\cong\mathbb{Z}/6\mathbb{Z}$
is generated by $[\mathfrak{q}_{1}]$ and $[\mathfrak{q}_{2}]$. Finally, the order $\mathcal{O}:=\mathbb{Z}+7{\tilde{\mathcal{O}}}$ also has conductor $7{\tilde{\mathcal{O}}}$ but index $7^{4}$ in ${\tilde{\mathcal{O}}}$. By Lemma 3.4 it only has a single prime ideal $\mathfrak{p}$ over $7$. Theorem 2.9 and the fact that $\mathcal{O}/\mathfrak{p}\cong\mathbb{F}_{7}$ show that $\operatorname{Chow}(\mathcal{O})$ must be trivial.
@@ 5. A maximality condition
As in Section 2, let $\mathfrak{p}_{1},\dots,\mathfrak{p}_{r}$ be the non-invertible maximal ideals of $\mathcal{O}$ and let $\mathfrak{P}_{i,1},\dots,\mathfrak{P}_{i,r_{i}}$ be the maximal ideals of ${\tilde{\mathcal{O}}}$ over $\mathfrak{p}_{i}$.
@@
We first show that (1) implies (2). So suppose $\mathcal{I}(\mathcal{O})\to\mathcal{I}({\tilde{\mathcal{O}}})$ is injective. If $u\in{\tilde{\mathcal{O}}}^{*}$, then $u\mathcal{O}$ is a principal ideal that is in the kernel of $\mathcal{I}(\mathcal{O})\to\mathcal{I}({\tilde{\mathcal{O}}})$. Hence $u\in\mathcal{O}^{*}$ and therefore $\mathcal{O}^{*}={\tilde{\mathcal{O}}}^{*}$. Suppose now $[\mathfrak{a}]$ is in the kernel of $\operatorname{Pic}(\mathcal{O})\to\operatorname{Pic}({\tilde{\mathcal{O}}})$. Then $\mathfrak{a}{\tilde{\mathcal{O}}}=\alpha{\tilde{\mathcal{O}}}$ for some $\alpha\in K^{*}$. Hence $\alpha^{-1}\mathfrak{a}$ is in the kernel of $\mathcal{I}(\mathcal{O})\to\mathcal{I}({\tilde{\mathcal{O}}})$. This implies $\alpha^{-1}\mathfrak{a}=\mathcal{O}$. Hence $[\mathfrak{a}]=[\mathcal{O}]$.
@@
see for example [8, Proposition 12.9]. The Chinese Remainder Theorem shows that (5) implies (4) which clearly implies (3). So it remains to show that (3) implies (5). For $x\in{\tilde{\mathcal{O}}}$ and $\pi\in\prod_{i=1}^{r}\prod_{j=1}^{r_{i}}\mathfrak{P}_{i,j}$ the ideal $(1+\pi x){\tilde{\mathcal{O}}}$ is coprime to $\mathcal{F}$ and thus
@@
Theorem 5.4 below is [9, Theorem 3.1]. Since the proof presented there relies on the wrong statement [9, Theorem 2.1], we decided to give a proof here. We start by recalling a well known fact.
@@
Suppose $\operatorname{div}_{\mathcal{O}}$ is injective and ${\tilde{\mathcal{O}}}\neq\mathcal{O}$. Then $\mathcal{O}^{*}={\tilde{\mathcal{O}}}^{*}$ by Corollary 2.2. Remark 2.7 shows that $\operatorname{div}_{\mathcal{O}}$ is the composition of the maps
@@
Hence $\mathcal{I}(\mathcal{O})\to\mathcal{I}({\tilde{\mathcal{O}}})$ must be injective and thus an isomorphism by Lemma 5.3. Thus $f_{*}$ must also be injective. Corollary 2.8 shows that $r_{i}=1$ for all $i$. But this contradicts Theorem 5.1. ∎
@@
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.

### v3

@@ 4. Further examples
The maximal order ${\tilde{\mathcal{O}}}$ of $K:=\mathbb{Q}(\sqrt{-3},\sqrt{13})$ has class number $2$. Lemma 3.1 shows that the order $\mathcal{O}=\mathbb{Z}+2{\tilde{\mathcal{O}}}$ has a unique maximal ideal $\mathfrak{p}$ over $2$. Then $2{\tilde{\mathcal{O}}}=\mathfrak{P}\mathfrak{Q}$ with maximal ideals $\mathfrak{P}$ and $\mathfrak{Q}$ which both generate the class group of ${\tilde{\mathcal{O}}}$. Since $\mathfrak{P}$ and $\mathfrak{Q}$ have degree $2$ over $\mathfrak{p}$, the local Chow group $\operatorname{Chow}(\mathcal{O}_{\mathfrak{p}})$ has order $2$. Theorem 2.6 shows
@@
splits into two maximal ideals with $[{\tilde{\mathcal{O}}}:\mathfrak{P}_{1}]=7^{2}$ and $[{\tilde{\mathcal{O}}}:\mathfrak{P}_{2}]=7^{3}$. The order
@@
has conductor $\mathfrak{P}_{1}$ and index $7$ in ${\tilde{\mathcal{O}}}$. The ideal $\mathfrak{P}_{1}$ is the unique non-invertible maximal ideal of $\mathcal{O}_{1}$ and $\mathcal{O}_{1}/\mathfrak{P}_{1}\cong\mathbb{F}_{7}$. Thus Theorem 2.6 shows that
@@
has conductor $\mathfrak{P}_{2}$ and index $7^{2}$ in ${\tilde{\mathcal{O}}}$. Again, $\mathfrak{P}_{2}$ is the unique non-invertible maximal ideal of $\mathcal{O}_{2}$ and $\mathcal{O}_{2}/\mathfrak{P}_{2}\cong\mathbb{F}_{7}$ implies that
@@
is generated by $[\mathfrak{P}_{2}]$. The intersection $\mathcal{O}^{\prime}:=\mathcal{O}_{1}\cap\mathcal{O}_{2}$ has conductor $7{\tilde{\mathcal{O}}}$ and index $7^{3}$ in ${\tilde{\mathcal{O}}}$. It has two different maximal ideals over $7$, namely $\mathfrak{p}_{i}:=\mathcal{O}^{\prime}\cap\mathfrak{P}_{i}$ for $i=1,2$. Theorem 2.6 shows that
$\operatorname{Chow}(\mathcal{O}^{\prime})\cong\operatorname{Chow}(\mathcal{O}^{\prime}_{\mathfrak{p}_{1}})\times\operatorname{Chow}(\mathcal{O}^{\prime}_{\mathfrak{p}_{1}})\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/3\mathbb{Z}\cong\mathbb{Z}/6\mathbb{Z}$
is generated by $[\mathfrak{p}_{1}]$ and $[\mathfrak{p}_{2}]$. Finally, the order $\mathcal{O}:=\mathbb{Z}+7{\tilde{\mathcal{O}}}$ also has conductor $7{\tilde{\mathcal{O}}}$ but index $7^{4}$ in ${\tilde{\mathcal{O}}}$. By Lemma 3.1 it only has a single maximal ideal $\mathfrak{p}$ over $7$. Theorem 2.6 and the fact that $\mathcal{O}/\mathfrak{p}\cong\mathbb{F}_{7}$ show that $\operatorname{Chow}(\mathcal{O})$ must be trivial.
@@ 5. A maximality condition
As in Section 2, let $\mathfrak{p}_{1},\dots,\mathfrak{p}_{r}$ be the non-invertible maximal ideals of $\mathcal{O}$ and let $\mathfrak{P}_{i,1},\dots,\mathfrak{P}_{i,r_{i}}$ be the maximal ideals of ${\tilde{\mathcal{O}}}$ over $\mathfrak{p}_{i}$. Further let $\mathcal{F}$ be the conductor of $\mathcal{O}$ in ${\tilde{\mathcal{O}}}$.
@@
We first show that (1) implies (2). So suppose $\mathcal{I}(\mathcal{O})\to\mathcal{I}({\tilde{\mathcal{O}}})$ is injective. If $u\in{\tilde{\mathcal{O}}}^{*}$, then the fractional principal ideal $u\mathcal{O}$ is in the kernel of $\mathcal{I}(\mathcal{O})\to\mathcal{I}({\tilde{\mathcal{O}}})$. Hence $u\in\mathcal{O}^{*}$ and therefore $\mathcal{O}^{*}={\tilde{\mathcal{O}}}^{*}$. Suppose now $[\mathfrak{a}]$ is in the kernel of $\operatorname{Pic}(\mathcal{O})\to\operatorname{Pic}({\tilde{\mathcal{O}}})$. Then $\mathfrak{a}{\tilde{\mathcal{O}}}=\alpha{\tilde{\mathcal{O}}}$ for some $\alpha\in K^{*}$. Hence $\alpha^{-1}\mathfrak{a}$ is in the kernel of $\mathcal{I}(\mathcal{O})\to\mathcal{I}({\tilde{\mathcal{O}}})$. This implies $\alpha^{-1}\mathfrak{a}=\mathcal{O}$. Hence $[\mathfrak{a}]=[\mathcal{O}]$.
@@
see for example [9, Proposition 12.9]. The Chinese Remainder Theorem shows that (5) implies (4) which clearly implies (3). So it remains to show that (3) implies (5). For $x\in{\tilde{\mathcal{O}}}$ and $\pi\in\prod_{i=1}^{r}\prod_{j=1}^{r_{i}}\mathfrak{P}_{i,j}$ the ideal $(1+\pi x){\tilde{\mathcal{O}}}$ is coprime to $\mathcal{F}$ and thus
@@
Theorem 5.4 below is [11, Theorem 3.1]. Since the proof presented there relies on the wrong statement [11, Theorem 2.1], we decided to give a proof here. We start by recalling a well known fact.
@@
Suppose $\operatorname{div}_{\mathcal{O}}$ is injective. Then $\mathcal{O}^{*}={\tilde{\mathcal{O}}}^{*}$ by Corollary 2.3. Equation (2.1) shows that $\operatorname{div}_{\mathcal{O}}$ is the composition of the maps
@@
Hence $\mathcal{I}(\mathcal{O})\to\mathcal{I}({\tilde{\mathcal{O}}})$ must be injective and thus an isomorphism by Lemma 5.3. Thus $f_{*}$ must also be injective. Lemma 2.5 shows that $r_{i}=1$ for all $i$. But this contradicts Theorem 5.1 if ${\tilde{\mathcal{O}}}\neq\mathcal{O}$. ∎
@@
Data sharing not applicable to this article as no datasets were generated or analyzed during the current study.
Declarations
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 491392403 – TRR 358.
The authors have no financial or proprietary interests in any material discussed in this article.